From nch@roe.ac.uk Fri Mar 9 15:07:53 2012 Date: Fri, 9 Mar 2012 15:07:51 +0000 From: Nigel Hambly To: Eckhard Sutorius , Nicholas Cross , Rob Blake , Mike Read , Ross Collins , Clive Davenhall , Mark Holliman Cc: Jim Emerson , Lorenzo Rimoldini , Stelios Voutsinas , Keith Noddle , Tom Shanks , Norman Gray , Mike Irwin , Andy Lawrence at ROE , Tom Kerr , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archives team meeting minutes, 9th March 2012 Present: ETWS, NJC, ACD, MAR, MSH, RSC, RGM, NCH, STV Apologies: JPE, AL, KTN, NG, RPB, C37, 89X DoNM: 10am Friday March 23rd 2012 in the VH (MAR in the chair) Actions discharged this time: ----------------------------- ACTION: NJC to contact Matt Jarvis and Will Sutherland to try and finalise the VIDEO and VIKING public release contents. Discharged; VIDEO oked but VIKING GAMA fields still uncertain Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB to roll out latest curation client / DB server share configuration for routine ops use. Continues; sahure has been done so far ACTION: ETWS to sort out common SQL function definition script(s) and ensure schema browser parser works correctly with those. An update to the schema browser parser is still required; ETWS noted that the schema parser needs a complete redesign to cope with the fragmented VISTA survey release policy. ACTION: MAR to update VSA pages with details of public survey releases. Done for the VHS; others will need done as they appear. Actions carried forward from 24/02/12 meeting --------------------------------------------- ACTION: RPB to set up WSA mirror/sync for flat file DBs using MS SQL replication. Continues; a list of tables to publish was agreed. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one has low priority so keeps getting trumped by more important things. ACTION: RPB to complete the LTO-4 tape backups of VVV and SSA DBs and SSS flat file datasets. Continues; a linux client has been installed and the SSS backed up. RPB noted earlier in the week that he's not at all confident that the backups of /scats and /scatn have worked as the verify fell over, so he will try redoing in smaller chunks. Specific points and new actions: -------------------------------- Project management: RGM noted that the WFAU grant renewal submission will now be in June. Hardware resourcing is still not clear; the old tried-and-tested method in operation up until now is no longer possible owing to restrictions on the Research Council's capital spend budgets). Some unminutable suggestions for workarounds were discussed, and some unprintable expletives uttered as to the craziness of the situation. WFCAM, VISTA and VST updates: No news to report this time. Comments and issues arising from CASU minutes: The team noted the minutes of the meeting of 22nd February. With regard to the heads-up over camera #3 snafus, ETWS notes that no problems arose with the missing detector #3 on 20120108 and 09 from the transfer/ingests done for January. Networking: ETWS noted that the latest batch of WSA data is all safely transfered and available for flat-file access; there were some issues over duplication of the data on disks at the CASU end but this has been sorted out now. WSA/VSA/OSA/GES Operations: NB: GES = Gaia-ESO Survey ETWS reported: -- Transferred the WSA January 2012 data and ingested the metadata. -- Rewriting Browser parser for VISTA releases in the VSA -- Worked on updates of the C++ code for the OSA. Apart from some schema quirks ingests work fine. -- Finalised draft browser pages with ACD for the GES ACD and NCH noted that the GES schema draft and a draft consortium message have been sent to IoA for feedabck, and that a request has been made for some example data to help populate table fields. NCH and ACD have also started work on the materials required for the Survey Implementation Plan, including a proto-ICD for the spectroscopic data. Hardware and Systems: MSH noted that the most recent hardware acquisitions (new DB and new processing nodes) will be arriving next week. Some disk juggling will then ensue to yield further storage expansion by 30TB and 50TB. A server will be configured specially for some MonetDB (column-oriented storage RDBMS) testing, and a small batch of 4 512GB SSDs has been acquired to do some testing with those. Software: NJC and RSC noted the usual SW maintenance of minor bug fixes and modifications. Survey Data Release: VISTA public release update: VVV: still processing synoptic products; ETA late March / early April VMC: Maria-Rosa has agreed to a public release to coincide with NAM. The database has been built from an existing consortium release and in line with the restrictions imposed by the PI and ESO. RSC noted that this leads to some inconsistencies within the database when some data-not-to-be-released are expunged, leaving any links and (for example) seaming information in overlapping regions broken. After some discussion, it was agreed to simply document such issues with the release webdocs rather than try to fix up. VIDEO: Contents now agreed with PI and database built; currently being tested by PI's postdoc. VIKING: Still waiting to hear back from Will over field selection/ deprecations, in particular the specification of which GAMA fields to be included. VHS: Following Richard McMahon's approval DR1 has been released, as noted in the last minutes. WSA: We anticipate rapid progress on GPS DR9 now that the calibration issues have been sorted out, and that the survey PI has hired a postdoc to help with eyeball QC. RGM asked about how the putative UHS will be served from the WSA, and there were some discussions but any proprietary issues need to be cleared up first, since we need to know if any of the data, or any of the extant UKIDSS data that is incorporated, are subject to proprietary rights. Non-survey Data Release: Nothing to report this time. Astrogrid deployment & Data Analysis services: NCH asked about the progress on UCD1+ and if there were some tests available in client-side software that presently don't work and that should spark in to life if we use the new standard. ETWS pointed out that newly parsed UCDs produced for SDSS-DR8 can be compared against DR7 which does have them. MSH noted that he has been given the go-ahead to migrate the Astrogrid servers and core services north from Leicester to maintain them in Edinburgh. Miscellaneous: VHS-UltraCoolDwarfs telecon takes place next Tuesday at 15:00 GMT. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. ============================================================= -- Scanned by iCritical.